OpenStudy (anonymous):

3^x = (1/9)^3 What is the value of x?

OpenStudy (akshay_budhkar):

1/9 = 3 ^-2 so 3^x= 3^-2*3 therefore x=-6

OpenStudy (anonymous):

alright thanks what about 4^x = 1

OpenStudy (akshay_budhkar):

4 ^x = 4^0 ( as 4^0=1) so x=0

OpenStudy (anonymous):

thanks and last one please x^-3 = 1/1000

OpenStudy (akshay_budhkar):

LOL ok

OpenStudy (anonymous):

Thanks a lot mate

OpenStudy (akshay_budhkar):

\[x^{-3} = 1/ x^3\] and \[1/1000 = 1/ 10 ^3\] so x= 10

OpenStudy (anonymous):

Thanks!

OpenStudy (akshay_budhkar):

yw :D